OpenStudy (anonymous):

what are the x-intercepts for f(x)= x^2-6x+3 ? please help!!!

OpenStudy (anonymous):

\[ 0 = x^2-6x+3= (x-3)^2-6 \] So the intercepts are \[ 3\pm\sqrt{6} \]